The decomposition of NO2(g) : 2NO2(g) ----->2NO(g) +O(g) is a second order reaction and the rate constant is 0.755L/(mol-s). In an experiment, the initial concentration of NO2(g) was 0.00650 mol/L. (a)What is the concentration of NO(g) after 125 seconds have elapsed? (b) How many seconds will it take for the concentration of NO2(g) to drop to 0.00100 mol/L?


1
Expert's answer
2021-02-24T03:38:10-0500

(a) For second order reaction:

"k = \frac{1}{t}[\frac{1}{A_t} -\frac{1}{A_0}] \\

0.755 = \frac{1}{125}[\frac{1}{A_t} -\frac{1}{0.0065}] \\

94.375 = \frac{1}{A_t} -\frac{1}{0.0065} \\

248.221 = \frac{1}{A_t} \\

A_t = 0.00402 \;M"

"(b) \; 0.755 = \frac{1}{t}[\frac{1}{0.00100} -\frac{1}{0.0065}] \\

0.755 = \frac{1}{t}[1000 – 153.84] \\

0.755 = \frac{846.16}{t} \\

t = 1120 \;s"